# TimeCopilot

### v1

@@ 2.2 TimeCopilot Forecaster
The TimeCopilot Forecaster executes the models proposed by the agent. It provides access to the largest unified hub of TSFMs under a single API  [^2] A complete and regularly updated list of models is available at https://timecopilot.dev/model_hub/, alongside statistical, machine learning, and deep learning methods. This hub removes the need for users to manage conflicting dependencies or learn multiple fragmented interfaces. The current set of supported TSFMs includes Chronos (Ansari et al., 2024), Moirai (Woo et al., 2024), Sundial (Liu et al., 2025b), TabPFN (Hoo et al., 2025), TiRex (Auer et al., 2025), TimeGPT (Garza et al., 2023a), TimesFM (Das et al., 2023), and Toto (Cohen et al., 2024).
@@ Acknowledgments
We would like to express our gratitude to Martin Juárez and Salomón Márquez for their helpful contributions and support.

### v2

@@ 2.2 TimeCopilot Forecaster
The TimeCopilot Forecaster executes the models proposed by the agent. It provides access to the largest unified hub of TSFMs under a single API  [^2] A complete and regularly updated list of models is available at https://timecopilot.dev/model-hub/, alongside statistical, machine learning, and deep learning methods. This hub removes the need for users to manage conflicting dependencies or learn multiple fragmented interfaces. The current set of supported TSFMs includes Chronos (Ansari et al., 2024), Moirai (Woo et al., 2024), Sundial (Liu et al., 2025b), TabPFN (Hoo et al., 2025), TiRex (Auer et al., 2025), TimeGPT (Garza et al., 2023a), TimesFM (Das et al., 2023), and Toto (Cohen et al., 2024).
@@ Acknowledgments
We would like to express our gratitude to Martin Juárez, Rodrigo Mendoza Smith, and Salomón Márquez for their helpful contributions and support.